OpenStudy (anonymous):

what is the derivative of (1+h)^ln(1+h)

myininaya (myininaya):

let y=(1+h)^(ln(1+h)) lny=ln(1+h)^(ln(1+h)) lny=ln(1+h)*ln(1+h) lny=[ln(1+h)]^2 take derivative of both sides y'/y=2ln(1+h)*1/(1+h)

myininaya (myininaya):

so y'={2yln(1+h)}/(1+h) y'=[(1+h)^(ln(1+h))*ln(1+h)]/(1+h)

myininaya (myininaya):

times 2